OpenStudy (anonymous):

Hello :) I was just wondering if I solved this problem correctly? 2n=8(n-6) n is six right?

OpenStudy (saifoo.khan):

hi

OpenStudy (saifoo.khan):

2n = 8n - 48 6n = 48 n = 48/6 n =

OpenStudy (akshay_budhkar):

thanks saif

OpenStudy (saifoo.khan):

uw

OpenStudy (anonymous):

Oh....ok. Thank you for showing my mistake :) I appreciate it.

OpenStudy (saifoo.khan):

np! :)

OpenStudy (anonymous):

So the answer to n is eight?

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

Alrighty. Thanks

OpenStudy (saifoo.khan):

welks.